OpenStudy (faman39):

Matching correct pair of equivalent ratios. 3 to 5 =20% 4/16 = 9 to10 36:40 =3:12 8/40 =12/20

OpenStudy (wasiqss):

36.40 =9/`10 4/16=3:12 8/40=20% 3/5=12/20

Directrix (directrix):

None of these statements are true. Perhaps I should reread the question because I don't see anything to match.

razor99 (razor99):

awooooooooooo

OpenStudy (wasiqss):

hahaah directrix you need a shock treatmet :D

razor99 (razor99):

hey i did not deserve it

razor99 (razor99):

@faman

razor99 (razor99):

hey chris

razor99 (razor99):

i think directrix is correct

OpenStudy (callisto):

At first glance, you'll reject all the options because you think the numbers doesn't match. On second thought, you should look at the question: 'Matching correct pair of equivalent ratios'. It actually asks you to match, not determine which is correct!

Directrix (directrix):

How the Question Should be Phrased ---------- Given the numbered four ratios on the left, write the letter of the correspondending ratio to which it is equivalent (if any): 1. 3 to 5 a) 20% 2. 4/16 b)9 to10 3. 36:40 c)3:12 4. 8/40 d) 12/20

OpenStudy (callisto):

for the right column, just do the same thing as i did for the left column. It would be quite time-consuming, but once you get used to these conversion, you don't have to make such kind of tables

OpenStudy (callisto):

Sorry i attached the wrong file :( This is the correct one